During the Drop Test (IEC 61010), how far above the wooden board on the floor do you need to position the DUT (device under test), in inches?

A.

40.0

B.

39.0

C.

39.4

D.

37.6

E.

38.2

A motor has a 3000 VAC nameplate rating. What should be the value of its DC voltage proof test after service?

A.

7000 VDC

B.

6720 VDC

C.

8960 VDC

D.

4200 VDC

E.

6000 VDC

F.

3840 VDC

G.

5600 VDC

H.

3000 VAC

I.

3000 VDC

J.

2880 VDC
Answer #1

(1) Drop Test standard:

DUT Weight (lb) Drop Height (inches)
1-20 30
21-40 24
41-60 18
61-100 12
101-150 8

So to know the exacy height, you should mention the weight of DUT.

(2) Standard applied DC voltage for proof test is given by

                                 V_{applied,DC} = \left(2*V_{rated,AC} + 1000 \right )\hspace{2mm}V

Here motor rating is 3000V, so

                               V_{applied,DC} = \left(2*3000 + 1000 \right )\hspace{2mm}V = 7000\hspace{2mm}V

Correct option is (a).